# Keeping Score: A Quantitative Analysis of How the CHI Community Appreciates Its Milestones

### v2

@@ 4.3. Does the CHI Community show a Preference for its Milestone Authors? (RQ4)
The CHI community is relatively large, with over a thousand authors attending the CHI Conference. While attended by many first-timers, some authors at CHI are repeat authors (Kaye, 2009). The Matthew effect in science (Merton, 1968) describes the tendency of successful authors to be cited more often. This accumulated advantage may also be present in the milestones at CHI, and we expect many milestones authors have authored more than one milestone.
@@
[…]
A bar chart depicting the number papers that authors have (co-)authored among the top-300 most cited milestones at CHI. While the majority of authors have only written one of the top-300 milestones, about 20% of the authors have (co-)authored multiple of the top-300 milestones.
To investigate this, we identify the authors of the 300 milestones most cited in the CHI community (over all proceedings years) and count the number (see Figure 7) and cumulative percentage (see Figure 8) of milestones these authors have (co-)authored. We find that while the majority ($n=615$, 80.3%) of the milestone authors have only authored one of the top-300 milestones, 151 authors (19.7% of the top-300 milestone authors) have (co-)authored more than one of the top-300 milestones. The 50 authors who have (co-)authored three or more of CHI’s 300 most cited milestones are listed in Figure 8. Chief among them are Phoebe Sengers, Paul Dourish, and Jacob O. Wobbrock who each have (co-)authored eight of the CHI community’s 300 most cited milestones. This is followed by John Zimmerman (7 milestones), Jodi L. Forlizzi (6 milestones), and others (see Figure 8). Together, the 50 authors have (co-)authored about 20% of the 300 milestones most cited by the CHI community (see secondary axis in Figure 8).
Figure 8. Fifty authors have published three or more of the 300 milestones most cited in the CHI community (teal bars and lower x-axis). The 50 authors have (co-)authored about 20% of the 300 milestones (red line and top x-axis). The milestones publication years are listed in parentheses.
[…]
Authors having authored papers three or more milestones in the 300 papers most-cited by the CHI community. The 50 authors have (co-)authored about 20% of the 300 milestones (red line and top x-axis).
@@ 5. Discussion
The CHI community continues to grow in size and impact. In this paper, we presented meta-research on how the community operates as a scientific field. The way how we cite our milestones is not only an intellectually satisfying exercise, it also concretely explains what kind of work goes appreciated at large in the community.
@@ 5.1. The CHI Community’s Milestones
This contrasts with some of the contribution types in HCI that are not represented in the list of top-cited milestones. In particular, milestones are not being cited for contributing prototypes and systems, which is one very popular contribution type at CHI. Spurred by the availability of large language models, it is now easier than ever to implement prototypes in HCI and write a paper about them. We can observe that some authors rush from one quickly implemented HCI study to the next (Noorden and Perkel, 2023). The methodological acceleration in HCI (Thomas et al., 2024) shifts our attention away from things that may have a long-lasting impact in HCI in favor of quick results. Some researchers have raised concerns about the quick-win approach found in some HCI research, calling it concerning and detrimental to the field of HCI (Arawjo, 2024). HCI has a strong case of “shiny object syndrome”, where short-lasting projects are prioritized over time-persisting milestones. We believe it is not likely that the increasingly popular “thin-wrapper” implementations with language models will become part of the set of milestone contributions in HCI. It is important for us to not forget the past milestones and their contributions, as they inform our methodologies and values in HCI.
@@
The “sweet spot” of relevance for papers cited at CHI is 5–15 years back in time. The 5–15 year citation range likely reflects a balance between novelty, relevance, accessibility, and consolidation of knowledge. Research within this range has typically undergone sufficient validation and follow-up, making it both trustworthy and broadly applicable, while remaining relevant to current advances. This time frame allows foundational ideas to influence new work, often through citation cascades that amplify visibility and integration into broader academic and practical contexts. Ideas from this period are also more likely to have been incorporated into curricula and practice, making them accessible and influential to a wide audience. Similar patterns are seen in other disciplines, such as physics and medicine, where impactful research takes years to be validated, cited, and adopted into subsequent innovations. Conversely, older work may become less cited as it is superseded by newer paradigms or rendered less accessible. This highlights the dynamic interplay between innovation, validation, and knowledge dissemination in shaping research impact.
@@
These trends are indicative of a community at a crossroads between remembrance and forgetfulness, a community that while honoring and valuing past contributions, is gravitating towards more recent work. Whether this is a problem or natural evolution is, of course, subjective. These issues are known and have, of course, also been grappled with earlier, e.g. through the CHI2030 Vision Task Force (2018–2019) that sought to address the future of the CHI Conference.
@@ 5.2. Matthew Effect and Authoring at CHI
We demonstrated how certain prolific authors contribute a lot to the community milestones. The 50 most highly-cited authors have (co-)authored about 20% of the top-300 milestones cited at CHI, as depicted in Figure 8. Reasons for this concentration of milestone authorship may vary. Obviously, the work itself by these authors is of high-quality and constitutes advances worth remembering in new and upcoming research. And some preferential attachment is normal in a research community. On the other hand, some of the citation patterns might be explained through other forms of influence: large researcher networks, access to labs with plenty of funding and, thus, future researchers citing the past work from the same lab, etc. These are all factors that we should, in all fairness, recognize. It may be rather difficult to attain a milestone status for one’s work, if it originates from an underdog position and from a relatively unknown laboratory. In any case, this concentration of milestone authorship suggests an influential core within the CHI community.
What does this mean for authors at CHI? The cumulative advantage of existing milestones makes it difficult for emerging authors to publish new milestones. While it may arguably be easier for authors to contribute an impactful milestone in one of CHI’s many sub-communities, authoring a milestone that concerns the whole community is hard, as new milestones compete with the CHI community’s attention and established milestones. CHI’s attachment to its one super milestone is the best example for this accumulated advantage. The accumulated advantage and focused attention of the community on some highly performing milestones, we argue, creates barriers to entry for emerging milestones and their authors. One could speculate on whether the CHI community is saturated with milestones, which would further contribute to barriers of entry. Braun and Clarke’s work, for instance, seems to receive fewer citations (relative to all citations in the CHI Proceedings) in recent years, which could indicate that the CHI community has reached saturation with this important research method. This suggests that the CHI community may be at a point where established milestones overshadow the emergence of new, impactful contributions. In the following case study on CHI’s super milestone, we explore how this influential milestone navigated these challenges, setting a precedent for how future milestones might overcome these barriers.
@@ 5.3. Case Study on CHI’s Super Milestone
However, Braun and Clarke have more recently acknowledged that many researchers apply their method not as it was originally intended (Braun et al., 2019; Braun and Clarke, 2019). The authors themselves now advocate for the ‘fading’ of their own milestone in favor of their more recent works on Reflexive Thematic Analysis that explain how their original idea was intended. In many cases, researchers only loosely apply the method as an umbrella for their (unstructured) qualitative investigations or they apply different versions of the approach, as delineated by Braun et al. (Braun et al., 2019). Kaltenhauser et al. found the same issue in their review of autoethnographies, noting that “existing methods are often executed and interpreted with much variety” (Kaltenhauser et al., 2024). This speaks to how researchers often (mis-)appropriate methods and milestones.
@@ 5.4. The Milestone Coefficient: A New Lens on Impact
The Milestone Coefficient is by no means intended as a tool for research assessment. The MC, as yet another bibliometric metric, will not help with addressing the current problems with bibliometrics and academia. Further, only very few authors will—by definition—become milestone authors. For this reason, the Milestone Coefficient would be ineffective for the purpose of evaluating the career progress of most researchers. Rather, our intention with the Milestone Coefficient is to provide a tool and quantitative means for identifying milestones, discussing, and comparing the performance of milestones, going beyond a discrete typology of milestones. We hope this coefficient will provoke a discussion on the works that have had the greatest impact at CHI and about what we value in the HCI research community.
@@ 5.5. Limitations and Future Work
Second, our work examines how the CHI community cites its milestones. It is, however, possible that the CHI community favors derivative works over milestones. Metaphorically, authors of derivative works stand on the shoulders of giants, but by becoming giants themselves, they may crush what is underneath them. One example of this is Fitts’ Law (Fitts, 1954). Fitts’ highly influential work was published in (1954) and inspired much research in HCI. However, Fitts’ success was shared by other authors who presented studies on Fitts Law. While we do not find strong evidence for other instances of this in our research, future work could investigate how newer milestones and derivative works displace attention to older milestones.
Third, we acknowledge that our method of identifying milestone papers and authors are imperfect. We uniquely identify papers by their title and publication year. We use regular expressions to extract the paper titles. While we use normalization (such as lower casing and removal of punctuation and white spaces), our simple string matching approach is limited in uniquely identifying papers. References in CHI papers are noisy and may contain human mistakes, such as errors and typos. Our method cannot address these human errors. In a similar vein, author name disambiguation is a common problem in bibliometric analyses that is impossible to solve completely (Bartneck and Hu, 2009; Kaye, 2009). For instance, authors may use different versions of their name, or they may legally change their name entirely. While we did our best to develop a set of regular expressions to capture different notations of author names, human errors can not be fully accounted for.
@@
Last, our methodology focuses on total citations over the entire ACM CHI conference. This methodology is not likely to identify milestones that have been popular for a very short period of time but since have faded completely from CHI’s collective memory. This would not allow the milestones to have accrued enough citations to be noticable in our methodology. We leave this and investigations of milestone impact in CHI’s sub-communities to future work.
@@ 6. Conclusion
We presented an analysis of how the CHI community is, over time, shifting its attention away from its past milestone papers. Forgetting curves provide graphical evidence that the relative attention of the CHI community to its milestone publications is decreasing over time. The attention span of the CHI community narrowly rests on milestone papers published in about ten year window. While this is a testament for the dynamism and renewal of research at CHI and in the field of HCI, it raises concerns about the value, sustainability, and meaningfulness of CHI research. The HCI field’s key contributions consist of methodological and conceptual contributions that move the field forward. However, it is unclear whether recent milestones can achieve the fame of past milestones in HCI.

### v3

@@ 4.3. Does the CHI Community show a Preference for its Milestone Authors? (RQ4)
The CHI community is relatively large, with over a thousand authors attending the CHI Conference. While attended by many first-timers, some authors at CHI are repeat authors (Kaye, 2009). The Matthew effect in science (Merton, 1968) describes the tendency of successful authors to be cited more often. This accumulated advantage may also be present in the milestones at CHI, and we expect many milestone authors have authored more than one milestone.
@@
To investigate this, we identify the authors of the 300 milestones most cited in the CHI community (over all proceedings years) and count the number (see Figure 7) and cumulative percentage (see Figure 8) of milestones these authors have (co-)authored. We find that while the majority ($n=615$, 80.3%) of the milestone authors have only authored one of the top-300 milestones, 151 authors (19.7% of the top-300 milestone authors) have (co-)authored more than one of the top-300 milestones. The 50 authors who have (co-)authored three or more of CHI’s 300 most cited milestones are listed in Figure 8. Chief among them are Phoebe Sengers, Paul Dourish, and Jacob O. Wobbrock who each have (co-)authored eight of the CHI community’s 300 most cited milestones. This is followed by John Zimmerman (7 milestones), Jodi L. Forlizzi (6 milestones), and others (see Figure 8). Together, the 50 authors have (co-)authored about 20% of the 300 milestones most cited by the CHI community (see the secondary axis on top of Figure 8).
Figure 8. Fifty authors have published three or more of the 300 milestones most cited in the CHI community (teal bars and lower x-axis). The 50 authors have (co-)authored 20% of the 300 milestones (red line and top x-axis). The milestones publication years are listed in parentheses.
@@ 5. Discussion
The CHI community continues to grow in size and impact. In this paper, we presented meta-research on how the community operates as a scientific field. Analyzing the way how we cite our milestones is not only an intellectually satisfying exercise, it also concretely explains what kind of work goes appreciated at large in the community.
@@ 5.1. The CHI Community’s Milestones
This contrasts with some of the contribution types in HCI that are not represented in the list of top-cited milestones. In particular, milestones are not being cited for contributing prototypes and systems, which is one very popular contribution type at CHI. Spurred by the availability of large language models, it is now easier than ever to implement prototypes in HCI or write a paper about them (Pang et al., 2025). We can observe that some authors rush from one quickly implemented HCI study to the next (Noorden and Perkel, 2023). The methodological acceleration in HCI (Thomas et al., 2024) shifts our attention away from things that may have a long-lasting impact in HCI in favor of quick results. Some researchers have raised concerns about the quick-win approach found in some HCI research, calling it concerning and detrimental to the field of HCI (Arawjo, 2024; Morris, 2024). HCI has a strong case of “shiny object syndrome”, where short-lasting projects are prioritized over time-persisting milestones. We believe it is not likely that the increasingly popular “thin-wrapper” implementations with language models will become part of the set of milestone contributions in HCI. It is important for us to not forget the past milestones and their contributions, as they inform our methodologies and values in HCI.
@@
The “sweet spot” of relevance for papers cited at CHI is 5–15 years back in time. This range likely reflects a balance between novelty, relevance, accessibility, and consolidation of knowledge. Research within this range has typically undergone sufficient validation and follow-up, making it both trustworthy and broadly applicable, while remaining relevant to current advances. This time frame allows foundational ideas to influence new work, often through citation cascades that amplify visibility and integration into broader academic and practical contexts. Ideas from this period are also more likely to have been incorporated into curricula and practice, making them accessible and influential to a wide audience. Similar patterns are seen in other disciplines, such as physics and medicine, where impactful research takes years to be validated, cited, and adopted into subsequent innovations. Conversely, older work may become less cited as it is superseded by newer paradigms or rendered less accessible. This highlights the dynamic interplay between innovation, validation, and knowledge dissemination in shaping research impact.
@@
These trends are indicative of a community at a crossroads between remembrance and forgetfulness, a community that while honoring and valuing past contributions, is gravitating towards more recent work. Whether this is a problem or natural evolution is, of course, up for debate. These issues are known and have, of course, also been grappled with earlier, for instance through the CHI2030 Vision Task Force (2018–2019) that sought to address the future of the CHI Conference.
@@ 5.2. Matthew Effect and Authoring at CHI
We demonstrated how certain prolific authors contribute a lot to the community milestones. The 50 most highly-cited authors have (co-)authored about 20% of the top-300 milestones cited at CHI, as depicted in Figure 8. There are numerous possible explanations for this concentration of prolific researchers. Obviously, the work itself by these authors is of high-quality and constitutes advances worth remembering in new and upcoming research in the field of HCI. And some preferential attachment is normal in a research community. On the other hand, some of the citation patterns might be explained through other forms of influence: large researcher networks, access to labs with plenty of funding and, thus, future researchers citing the past work from the same lab, etc. These are all factors that we should, in all fairness, recognize. It may be rather difficult to attain a milestone status for one’s work, if it originates from an underdog position and from a relatively unknown laboratory. In any case, we presented empirical evidence of a concentration of prolific milestone authorship in the CHI community.
What does this mean for authors at CHI? The cumulative advantage and prominence of existing milestones may make it difficult for emerging authors to publish new milestones. While it may arguably be easier for authors to contribute an impactful milestone in one of CHI’s many sub-communities, authoring a milestone that concerns the whole community is hard, as new milestones may compete with the CHI community’s attention and established milestones. CHI’s attachment to its one super milestone is the best example for this accumulated advantage. Is a grounded theory approach truly the best way to analyze (big) data in the twenty-first century? Some researchers have expressed their doubts and criticms about this (Deterding and Waters, 2021). The accumulated advantage and focused attention of the community on some highly performing milestones, we argue, may create barriers to entry for emerging milestones and their authors.
One could speculate on whether the CHI community is saturated with milestones, which would further contribute to barriers of entry. Braun and Clarke’s work, for instance, seems to receive fewer citations in recent years (relative to all citations in the CHI Proceedings), which could indicate that the CHI community has reached saturation with this research method. This may suggest that the CHI community is at a point where established milestones overshadow the emergence of new, impactful contributions. In the following case study on CHI’s super milestone, we explore how Braun and Clarke’s influential milestone navigated these challenges, setting a precedent for how future milestones might overcome these potential barriers.
@@ 5.3. Case Study on CHI’s Super Milestone
However, Braun and Clarke have more recently acknowledged that many researchers apply their method not as it was originally intended (Braun et al., 2019; Braun and Clarke, 2019). The authors themselves now advocate for the ‘fading’ of their own milestone in favor of their more recent works on Reflexive Thematic Analysis. In many cases, researchers only loosely apply the method as an umbrella for their (unstructured) qualitative investigations or they apply different versions of the approach, as delineated by Braun et al. (Braun et al., 2019). Kaltenhauser et al. found the same issue in their review of autoethnographies, noting that “existing methods are often executed and interpreted with much variety” (Kaltenhauser et al., 2024). This speaks to how researchers often (mis-)appropriate methods and milestones.
@@ 5.4. The Milestone Coefficient: A New Lens on Impact
The Milestone Coefficient is by no means intended as a tool for research assessment or a substitute to the h-index. The MC, as yet another bibliometric metric, will not help with addressing the current problems with bibliometrics and academia. Further, only very few authors will—by definition—become milestone authors. For this reason, the Milestone Coefficient would be ineffective for the purpose of evaluating the career progress of most researchers. Rather, our intention with the Milestone Coefficient is to provide a tool and quantitative means for identifying milestones, discussing, and comparing the performance of highly-cited milestones, going beyond a discrete typology of milestones. We hope this coefficient will provoke a discussion on the works that have had the greatest impact at CHI and about what we value in the HCI research community.
@@ 5.5. Limitations and Future Work
Second, our work examines how the CHI community cites its milestones. It is, however, possible that the CHI community favors derivative works over the original milestones. Metaphorically, authors of derivative works stand on the shoulders of giants, but by becoming giants themselves, they may crush what is underneath them. One example of this is Fitts’ Law (Fitts, 1954). Fitts’ highly influential work was published in (1954) and inspired much research in HCI. However, Fitts’ success was shared by other authors who presented studies on Fitts Law. While we do not find strong evidence for other instances of this in our research, future work could investigate how newer milestones and derivative works displace attention to older milestones.
Third, we acknowledge that our method of identifying milestone papers and authors is imperfect. We uniquely identify papers by their title and publication year. We use regular expressions to extract the paper titles. While we use normalization (such as lower casing and removal of punctuation and white spaces), our simple string matching approach is limited in uniquely identifying papers. References in CHI papers are noisy and may contain human mistakes, such as errors and typos. Our method cannot address these human errors. In a similar vein, author name disambiguation is a common problem in bibliometric analyses that is impossible to solve completely (Bartneck and Hu, 2009; Kaye, 2009). For instance, authors may use different versions of their name (e.g., see the first author in (Hosio et al., 2018b) and (Hosio et al., 2018a)), or they may legally change their name entirely. While we did our best to develop a set of regular expressions to capture different notations of author names, human errors can not be fully accounted for.
@@
Last, our methodology focuses on total citations over the entire ACM CHI Conference. This methodology is not likely to identify milestones that have been popular for a very short period of time but since have faded completely from CHI’s collective memory. This would not allow the milestones to have accrued enough citations to be noticeable in our methodology. We leave this and investigations of milestone impact in CHI’s sub-communities to future work.
@@ 6. Conclusion
We presented an analysis of how the CHI community is, over time, shifting its attention away from its past milestone papers. Forgetting curves provide graphical evidence that the relative attention of the CHI community to its milestone publications is decreasing over time. The attention span of the CHI community narrowly rests on milestone papers published in about a ten year window. While this is a testament for the dynamism and renewal of research at CHI and in the field of HCI, it raises concerns about the value, sustainability, and meaningfulness of CHI research. The HCI field’s key contributions consist of methodological and conceptual contributions that move the field forward. However, it is unclear whether recent milestones can achieve the fame of past milestones in HCI.